A study found that the mean amount of time cars spent in​ drive-throughs of a certain​ fast-food restaurant was

139.9

seconds. Assuming​ drive-through times are normally distributed with a standard deviation of

31

​seconds, complete parts​ (a) through​ (d) below.

Click here to view the standard normal distribution table (page 1).

LOADING...

Click here to view the standard normal distribution table (page 2).

LOADING...

​(a) What is the probability that a randomly selected car will get through the​ restaurant's drive-through in less than

88

​seconds?

The probability that a randomly selected car will get through the​ restaurant's drive-through in less than

88

seconds is

nothing

.

​(Round to four decimal places as​ needed.)
Answer #1

a) P(Mean less than 88)

= P(X < 88)

= P(z < (88 - 139.9)/31)

= P(z < -1.67)

= 0.0470